OpenStudy (anonymous):

condense the expression 1/2 log5 16-3log5 x+4log5 y

OpenStudy (anonymous):

hmmm ? I don't understand the equation.

OpenStudy (anonymous):

\[1/2 \log_{5} 16- 3\log_{5} x+ 4\log_{5} y\]

OpenStudy (asnaseer):

do you know all the log rules?

OpenStudy (asnaseer):

if not - consult this page: http://www.purplemath.com/modules/logrules.htm you should then be able to solve this

OpenStudy (anonymous):

yes i think i do

OpenStudy (asnaseer):

an example step would be:\[3\log_5x=\log_5x^3\]

OpenStudy (anonymous):

\[\frac 12 \log_5 16-3 \log_5 x+4 \log_5 y = \log_5 \frac{4 y^4}{x^3}\]

OpenStudy (asnaseer):

I guess FoolForMath has given away the answer - but try and do it yourself to convince yourself that this is the right answer :)

OpenStudy (anonymous):

Yes I guess I am wrong.